Hey,
I've tried solving the following 2 equations but finding it hard to complete. Are you able to help me please?
Answers either as integers, or as fractions in their simplest form
1) 13x + 4 = 5x + 20
2) 3 - 2(x - 3)= 5/2
Are you able to help me with them, be much apprenticed. Many thanks

Answer by ewatrrr(24785)

 
Hi
1) 8x = 16, x = 2
2)3 - 2x + 6 = 5/2
9 - 5/2 = 2x
18 - 5 = 4x
13 = 4x
13/4 = x